# The MUSE eXtremely Deep Field: Detections of circumgalactic Siāii*Ā emission at ā ā thanks: Based on observations made with ESO telescope at the La Silla Paranal Observatory under the large program 1101.A-0127

### v1

@@ Acknowledgements.
HK acknowledges support from the Japan Society for the Promotion of Science (JSPS) Overseas Research Fellowship (202160056) as well as JSPS Research Fellowships for Young Scientists (202300224, 23KJ2148). HK thanks Yuri Nagai, an academic assistant at NAOJ, for her wonderful support. VM acknowledges support from the Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO) grant 016.VIDI.189.162 (āODINā). AV and TG are supported by the SNF grant PP00P2 211023. T.N. acknowledges support from Australian Research Council Laureate Fellowship FL180100060. I.P. acknowledges funding by the European Research Council through ERC-AdG SPECMAP-CGM, GA 101020943. This work is based on observations taken by VLT, which is operated by European Southern Observatory. This research made use of Astropy666http://www.astropy.org, which is a community-developed core Python package for Astronomy (The Astropy Collaboration etĀ al. 2013, 2018), and other software and packages:MPDAF (Piqueras etĀ al. 2019), PHOTUTILS, Numpy (Harris etĀ al. 2020), Scipy (Virtanen etĀ al. 2020), and matplotlib (Hunter 2007).References
- Bacon etĀ al. (2010) Bacon, R., Accardo, M., Adjali, L., etĀ al. 2010, Proceedings of the SPIE, 7735, 773508, conference Name: Ground-based and Airborne Instrumentation for Astronomy III Place: eprint: arXiv:2211.16795 ADS Bibcode: 2010SPIE.7735E..08B
- Bacon etĀ al. (2023) Bacon, R., Brinchmann, J., Conseil, S., etĀ al. 2023, Astronomy & Astrophysics, 670, A4
- Bacon etĀ al. (2017) Bacon, R., Conseil, S., Mary, D., etĀ al. 2017, Astronomy & Astrophysics, 608, A1
- Berry etĀ al. (2012) Berry, M., Gawiser, E., Guaita, L., etĀ al. 2012, The Astrophysical Journal, 749, 4
- Blaizot etĀ al. (2023) Blaizot, J., Garel, T., Verhamme, A., etĀ al. 2023, Monthly Notices of the Royal Astronomical Society, 523, 3749, aDS Bibcode: 2023MNRAS.523.3749B
- Bordoloi etĀ al. (2023) Bordoloi, R., Simcoe, R.Ā A., Matthee, J., etĀ al. 2023, EIGER IV: The cool 10$^4$K circumgalactic environment of high-$z$ galaxies reveals remarkably efficient IGM enrichment, arXiv:2307.01273 [astro-ph]
- Bradley etĀ al. (2021) Bradley, L., SipÅcz, B., Robitaille, T., etĀ al. 2021, astropy/photutils: 1.3.0
- Burchett etĀ al. (2021) Burchett, J.Ā N., Rubin, K. H.Ā R., Prochaska, J.Ā X., etĀ al. 2021, The Astrophysical Journal, 909, 151
- Carniani etĀ al. (2023) Carniani, S., Venturi, G., Parlanti, E., etĀ al. 2023, JADES: The incidence rate and properties of galactic outflows in low-mass galaxies across 3 < z < 9, publication Title: arXiv e-prints ADS Bibcode: 2023arXiv230611801C
- Carr etĀ al. (2018) Carr, C., Scarlata, C., Panagia, N., & Henry, A. 2018, The Astrophysical Journal, 860, 143
- Chen etĀ al. (2014) Chen, H.-W., Gauthier, J.-R., Sharon, K., etĀ al. 2014, Monthly Notices of the Royal Astronomical Society, 438, 1435
- Chen etĀ al. (2020) Chen, H.-W., Zahedy, F.Ā S., Boettcher, E., etĀ al. 2020, Monthly Notices of the Royal Astronomical Society, 497, 498, publisher: OUP ADS Bibcode: 2020MNRAS.497..498C
- Chen etĀ al. (2021) Chen, Y., Steidel, C.Ā C., Erb, D.Ā K., etĀ al. 2021, Monthly Notices of the Royal Astronomical Society, 508, 19
- Claeyssens etĀ al. (2022) Claeyssens, A., Richard, J., Blaizot, J., etĀ al. 2022, Astronomy & Astrophysics, 666, A78
- Dessauges-Zavadsky etĀ al. (2010) Dessauges-Zavadsky, M., DāOdorico, S., Schaerer, D., etĀ al. 2010, Astronomy and Astrophysics, 510, A26, aDS Bibcode: 2010A&Aā¦510A..26D
- DāEugenio etĀ al. (2024) DāEugenio, F., Cameron, A.Ā J., Scholtz, J., etĀ al. 2024, JADES Data Release 3 ā NIRSpec/MSA spectroscopy for 4,000 galaxies in the GOODS fields, publication Title: arXiv e-prints ADS Bibcode: 2024arXiv240406531D
- Du etĀ al. (2018) Du, X., Shapley, A.Ā E., Reddy, N.Ā A., etĀ al. 2018, The Astrophysical Journal, 860, 75
- Dutta etĀ al. (2023) Dutta, R., Fossati, M., Fumagalli, M., etĀ al. 2023, Monthly Notices of the Royal Astronomical Society, 522, 535
@@
- Dutta etĀ al. (2024) Dutta, S., Muzahid, S., Schaye, J., etĀ al. 2024, Monthly Notices of the Royal Astronomical Society, 528, 3745, aDS Bibcode: 2024MNRAS.528.3745D
- Epinat etĀ al. (2018) Epinat, B., Contini, T., Finley, H., etĀ al. 2018, Astronomy & Astrophysics, 609, A40
- Erb etĀ al. (2023) Erb, D.Ā K., Li, Z., Steidel, C.Ā C., etĀ al. 2023, The Astrophysical Journal, 953, 118, aDS Bibcode: 2023ApJā¦953..118E
- Erb etĀ al. (2010) Erb, D.Ā K., Pettini, M., Shapley, A.Ā E., etĀ al. 2010, The Astrophysical Journal, 719, 1168
- Erb etĀ al. (2012) Erb, D.Ā K., Quider, A.Ā M., Henry, A.Ā L., & Martin, C.Ā L. 2012, The Astrophysical Journal, 759, 26
- Feltre etĀ al. (2020) Feltre, A., Maseda, M.Ā V., Bacon, R., etĀ al. 2020, Astronomy & Astrophysics, 641, A118
@@
- Ferland etĀ al. (1998) Ferland, G.Ā J., Korista, K.Ā T., Verner, D.Ā A., etĀ al. 1998, Publications of the Astronomical Society of the Pacific, 110, 761, publisher: The University of Chicago Press
- Finley et al. (2017) Finley, H., Bouché, N., Contini, T., et al. 2017, Astronomy & Astrophysics, 605, A118
- Fossati etĀ al. (2019) Fossati, M., Fumagalli, M., Lofthouse, E.Ā K., etĀ al. 2019, Monthly Notices of the Royal Astronomical Society, 490, 1451, publisher: OUP ADS Bibcode: 2019MNRAS.490.1451F
- France etĀ al. (2010) France, K., Nell, N., Green, J.Ā C., & Leitherer, C. 2010, The Astrophysical Journal, 722, L80
- Fujimoto etĀ al. (2019) Fujimoto, S., Ouchi, M., Ferrara, A., etĀ al. 2019, The Astrophysical Journal, 887, 107
- Gazagnes etĀ al. (2023) Gazagnes, S., Mauerhofer, V., Berg, D.Ā A., etĀ al. 2023, The Astrophysical Journal, 952, 164, aDS Bibcode: 2023ApJā¦952..164G
- Ginolfi et al. (2020) Ginolfi, M., Jones, G. C., Béthermin, M., et al. 2020, Astronomy & Astrophysics, 633, A90
- Greene etĀ al. (2022) Greene, J., Bezanson, R., Ouchi, M., Silverman, J., & Group, t. P. G. E.Ā W. 2022, The Prime Focus Spectrograph Galaxy Evolution Survey, arXiv:2206.14908 [astro-ph]
- Guo et al. (2023a) Guo, Y., Bacon, R., Bouché, N. F., et al. 2023a, Nature, 624, 53, arXiv:2312.05167 [astro-ph]
- Guo etĀ al. (2023b) Guo, Y., Bacon, R., Wisotzki, L., etĀ al. 2023b, Spatially-resolved Spectroscopic Analysis of Ly$\alpha$ Haloes: Radial Evolution of the Ly$\alpha$ Line Profile out to 60 kpc, arXiv:2309.06311 [astro-ph]
- Guo etĀ al. (2023c) Guo, Y., Bacon, R., Wisotzki, L., etĀ al. 2023c, Median Surface Brightness Profiles of Lyman-$\alpha$ Haloes in the MUSE Extremely Deep Field, arXiv:2309.05513 [astro-ph]
- Harris etĀ al. (2020) Harris, C.Ā R., Millman, K.Ā J., Van DerĀ Walt, S.Ā J., etĀ al. 2020, Nature, 585, 357
- Herenz etĀ al. (2017) Herenz, E.Ā C., Urrutia, T., Wisotzki, L., etĀ al. 2017, Astronomy & Astrophysics, 606, A12
- Hunter (2007) Hunter, J.Ā D. 2007, Computing in Science & Engineering, 9, 90
- Illingworth etĀ al. (2013) Illingworth, G.Ā D., Magee, D., Oesch, P.Ā A., etĀ al. 2013, The Astrophysical Journal Supplement Series, 209, 6
- Inami etĀ al. (2017) Inami, H., Bacon, R., Brinchmann, J., etĀ al. 2017, Astronomy & Astrophysics, 608, A2
- Jaskot & Oey (2014) Jaskot, A.Ā E. & Oey, M.Ā S. 2014, The Astrophysical Journal, 791, L19
- Johnson etĀ al. (2018) Johnson, S.Ā D., Chen, H.-W., Straka, L.Ā A., etĀ al. 2018, The Astrophysical Journal, 869, L1
- Kerutt etĀ al. (2022) Kerutt, J., Wisotzki, L., Verhamme, A., etĀ al. 2022, Astronomy & Astrophysics, 659, A183
- Kramarenko etĀ al. (2023) Kramarenko, I.Ā G., Kerutt, J., Verhamme, A., etĀ al. 2023, Monthly Notices of the Royal Astronomical Society, 527, 9853, arXiv:2305.07044 [astro-ph]
- Kusakabe etĀ al. (2022) Kusakabe, H., Verhamme, A., Blaizot, J., etĀ al. 2022, Astronomy & Astrophysics, 660, A44
- Leclercq etĀ al. (2020) Leclercq, F., Bacon, R., Verhamme, A., etĀ al. 2020, Astronomy & Astrophysics, 635, A82
- Leclercq etĀ al. (2017) Leclercq, F., Bacon, R., Wisotzki, L., etĀ al. 2017, Astronomy & Astrophysics, 608, A8
- Leclercq etĀ al. (2022) Leclercq, F., Verhamme, A., Epinat, B., etĀ al. 2022, Astronomy & Astrophysics, 663, A11, arXiv:2203.05614 [astro-ph]
- Lee etĀ al. (2014) Lee, K.-G., Hennawi, J.Ā F., Stark, C., etĀ al. 2014, The Astrophysical Journal, 795, L12
- Lehner etĀ al. (2016) Lehner, N., OāMeara, J.Ā M., Howk, J.Ā C., Prochaska, J.Ā X., & Fumagalli, M. 2016, The Astrophysical Journal, 833, 283
- Lofthouse etĀ al. (2020) Lofthouse, E.Ā K., Fumagalli, M., Fossati, M., etĀ al. 2020, Monthly Notices of the Royal Astronomical Society, 491, 2057, publisher: OUP ADS Bibcode: 2020MNRAS.491.2057L
- Lopez etĀ al. (2018) Lopez, S., Tejos, N., Ledoux, C., etĀ al. 2018, Nature, 554, 493, aDS Bibcode: 2018Natur.554..493L
- Luo etĀ al. (2017) Luo, B., Brandt, W.Ā N., Xue, Y.Ā Q., etĀ al. 2017, The Astrophysical Journal Supplement Series, 228, 2, publisher: IOP ADS Bibcode: 2017ApJS..228ā¦.2L
- Martin etĀ al. (2013) Martin, C.Ā L., Shapley, A.Ā E., Coil, A.Ā L., etĀ al. 2013, The Astrophysical Journal, 770, 41
- Maseda etĀ al. (2023) Maseda, M.Ā V., Lewis, Z., Matthee, J., etĀ al. 2023, JWST/NIRSpec Measurements of Extremely Low Metallicities in High Equivalent Width Lyman-$\alpha$ Emitters, arXiv:2304.08511 [astro-ph]
- Mauerhofer etĀ al. (2021) Mauerhofer, V., Verhamme, A., Blaizot, J., etĀ al. 2021, Astronomy & Astrophysics, 646, A80
- Meyer etĀ al. (2024) Meyer, R.Ā A., Oesch, P.Ā A., Giovinazzo, E., etĀ al. 2024, JWST FRESCO: a comprehensive census of H$\beta$+[OIII] emitters at 6.8, publication Title: arXiv e-prints ADS Bibcode: 2024arXiv240505111M
- Moffat (1969) Moffat, A. F.Ā J. 1969, Astronomy and Astrophysics, 3, 455, aDS Bibcode: 1969A&Aā¦..3..455M
- Momose etĀ al. (2014) Momose, R., Ouchi, M., Nakajima, K., etĀ al. 2014, Monthly Notices of the Royal Astronomical Society, 442, 110
- Morrissey (2018) Morrissey, P. 2018, The Astrophysical Journal
- Méndez-HernÔndez et al. (2022) Méndez-HernÔndez, H., Cassata, P., Ibar, E., et al. 2022, Astronomy & Astrophysics, 666, A56, arXiv:2206.08923 [astro-ph]
- Oke & Gunn (1983) Oke, J.Ā B. & Gunn, J.Ā E. 1983, The Astrophysical Journal, 266, 713, aDS Bibcode: 1983ApJā¦266..713O
@@
- Piqueras etĀ al. (2019) Piqueras, L., Conseil, S., Shepherd, M., etĀ al. 2019, Astronomical Data Analysis Software and Systems XXVI ASP Conference Series, 521, 545, conference Name: Astronomical Data Analysis Software and Systems XXVI ADS Bibcode: 2019ASPC..521..545P
- Planck Collaboration etĀ al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., etĀ al. 2020, Astronomy & Astrophysics, 641, A6
- Prochaska etĀ al. (2011a) Prochaska, J.Ā X., Kasen, D., & Rubin, K. 2011a, The Astrophysical Journal, 734, 24
- Prochaska etĀ al. (2011b) Prochaska, J.Ā X., Weiner, B., Chen, H.-W., Mulchaey, J., & Cooksey, K. 2011b, The Astrophysical Journal, 740, 91
- Péroux & Howk (2020) Péroux, C. & Howk, J. C. 2020, Annual Review of Astronomy and Astrophysics, 58, 363
- Rafelski etĀ al. (2015) Rafelski, M., Teplitz, H.Ā I., Gardner, J.Ā P., etĀ al. 2015, The Astronomical Journal, 150, 31
- Rakic etĀ al. (2012) Rakic, O., Schaye, J., Steidel, C.Ā C., & Rudie, G.Ā C. 2012, The Astrophysical Journal
- Richard etĀ al. (2019) Richard, J., Bacon, R., Blaizot, J., etĀ al. 2019, BlueMUSE: Project Overview and Science Cases, arXiv:1906.01657 [astro-ph]
- Richard etĀ al. (2021) Richard, J., Claeyssens, A., Lagattuta, D., etĀ al. 2021, Astronomy & Astrophysics, 646, A83
- Rubin etĀ al. (2018) Rubin, K. H.Ā R., OāMeara, J.Ā M., Cooksey, K.Ā L., etĀ al. 2018, The Astrophysical Journal, 859, 146
- Rubin et al. (2011) Rubin, K. H. R., Prochaska, J. X., Ménard, B., et al. 2011, The Astrophysical Journal, 728, 55
- Schroetter et al. (2021) Schroetter, I., Bouché, N. F., Zabl, J., et al. 2021, Monthly Notices of the Royal Astronomical Society, 506, 1355
- Shaban etĀ al. (2022) Shaban, A., Bordoloi, R., Chisholm, J., etĀ al. 2022, The Astrophysical Journal, 936, 77, aDS Bibcode: 2022ApJā¦936ā¦77S
- Shapley etĀ al. (2003) Shapley, A.Ā E., Steidel, C.Ā C., & Pettini, M. 2003, The Astrophysical Journal, 588
- Shull & van Steenberg (1982) Shull, J.Ā M. & van Steenberg, M. 1982, The Astrophysical Journal Supplement Series, 48, 95, publisher: IOP ADS Bibcode: 1982ApJSā¦48ā¦95S

### v2

@@ Acknowledgements.
We thank the anonymous referee for constructive comments and suggestions. We would like to express our gratitude to Charlotte Simmonds for useful comments and suggestions. HK acknowledges support from the Japan Society for the Promotion of Science (JSPS) Overseas Research Fellowship (202160056) as well as JSPS Research Fellowships for Young Scientists (202300224, 23KJ2148). HK thanks Yuri Nagai, an academic assistant at NAOJ, for her wonderful support. VM acknowledges support from the Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO) grant 016.VIDI.189.162 (’ODIN’). AV and TG are supported by the SNF grant PP00P2 211023. T.N. acknowledges support from Australian Research Council Laureate Fellowship FL180100060. I.P. acknowledges funding by the European Research Council through ERC-AdG SPECMAP-CGM, GA 101020943. This work is based on observations taken by VLT, which is operated by European Southern Observatory. This research made use of Astropy888http://www.astropy.org, which is a community-developed core Python package for Astronomy (The Astropy Collaboration et al. 2013, 2018), and other software and packages:MPDAF (Piqueras et al. 2019), PHOTUTILS, Numpy (Harris et al. 2020), Scipy (Virtanen et al. 2020), and matplotlib (Hunter 2007).References
- Bacon et al. (2010) Bacon, R., Accardo, M., Adjali, L., et al. 2010, Proceedings of the SPIE, 7735, 773508, arXiv eprints [arXiv:211.16795]
- Bacon et al. (2023) Bacon, R., Brinchmann, J., Conseil, S., et al. 2023, A&A, 670, A4
- Bacon et al. (2017) Bacon, R., Conseil, S., Mary, D., et al. 2017, A&A, 608, A1
- Banerjee et al. (2023) Banerjee, E., Muzahid, S., Schaye, J., Johnson, S. D., & Cantalupo, S. 2023, MNRAS, 524, 5148
- Beckett et al. (2024) Beckett, A., Rafelski, M., Revalski, M., et al. 2024, arXiv eprints [arXiv:2408.11914]
- Behroozi & Silk (2018) Behroozi, P. & Silk, J. 2018, Monthly Notices of the Royal Astronomical Society, 477, 5382
- Berry et al. (2012) Berry, M., Gawiser, E., Guaita, L., et al. 2012,ApJ, 749, 4
- Blaizot et al. (2023) Blaizot, J., Garel, T., Verhamme, A., et al. 2023, MNRAS, 523, 3749
- Bordoloi et al. (2024) Bordoloi, R., Simcoe, R. A., Matthee, J., et al. 2024, ApJ, 963, 28
- Bradley et al. (2021) Bradley, L., Sipőcz, B., Robitaille, T., et al. 2021, astropy/photutils: 1.3.0
- Burchett et al. (2021) Burchett, J. N., Rubin, K. H. R., Prochaska, J. X., et al. 2021, ApJ, 909, 151
- Carniani et al. (2024) Carniani, S., Venturi, G., Parlanti, E., et al. 2024, Astronomy and Astrophysics, 685, A99
- Carr et al. (2018) Carr, C., Scarlata, C., Panagia, N., & Henry, A. 2018, ApJ, 860, 143
- Chen et al. (2014) Chen, H.-W., Gauthier, J.-R., Sharon, K., et al. 2014, MNRAS, 438, 1435
- Chen et al. (2020) Chen, H.-W., Zahedy, F. S., Boettcher, E., et al. 2020, MNRAS, 497, 498
- Chen et al. (2021) Chen, Y., Steidel, C. C., Erb, D. K., et al. 2021, MNRAS, 508, 19
- Claeyssens et al. (2022) Claeyssens, A., Richard, J., Blaizot, J., et al. 2022, A&A, 666, A78
- Dessauges-Zavadsky et al. (2010) Dessauges-Zavadsky, M., D’Odorico, S., Schaerer, D., et al. 2010, A&A, 510, A26
- D’Eugenio et al. (2024) D’Eugenio, F., Cameron, A. J., Scholtz, J., et al. 2024, arXiv e-prints [arXiv: 2404.06531]
- Du et al. (2018) Du, X., Shapley, A. E., Reddy, N. A., et al. 2018,ApJ, 860, 75
- Dutta et al. (2023) Dutta, R., Fossati, M., Fumagalli, M., et al. 2023, MNRAS, 522, 535
@@
- Dutta et al. (2024) Dutta, S., Muzahid, S., Schaye, J., et al. 2024, Monthly Notices of the Royal Astronomical Society, 528, 3745
- Epinat et al. (2018) Epinat, B., Contini, T., Finley, H., et al. 2018, A&A, 609, A40
- Erb et al. (2023) Erb, D. K., Li, Z., Steidel, C. C., et al. 2023,ApJ, 953, 118
- Erb et al. (2010) Erb, D. K., Pettini, M., Shapley, A. E., et al. 2010, ApJ, 719, 1168
- Erb et al. (2012) Erb, D. K., Quider, A. M., Henry, A. L., & Martin, C. L. 2012, ApJ, 759, 26
- Feltre et al. (2020) Feltre, A., Maseda, M. V., Bacon, R., et al. 2020, A&A, 641, A118
@@
- Ferland et al. (1998) Ferland, G. J., Korista, K. T., Verner, D. A., et al. 1998, Publications of the Astronomical Society of the Pacific, 110, 761, publisher: The University of Chicago Press
- Finley et al. (2017) Finley, H., Bouché, N., Contini, T., et al. 2017, A&A, 605, A118
- Fossati et al. (2019) Fossati, M., Fumagalli, M., Lofthouse, E. K., et al. 2019, MNRAS, 490, 1451
- France et al. (2010) France, K., Nell, N., Green, J. C., & Leitherer, C. 2010, ApJ, 722, L80
- Fujimoto et al. (2019) Fujimoto, S., Ouchi, M., Ferrara, A., et al. 2019,ApJ, 887, 107
- Gazagnes et al. (2023) Gazagnes, S., Mauerhofer, V., Berg, D. A., et al. 2023, ApJ, 952, 164
- Ginolfi et al. (2020) Ginolfi, M., Jones, G. C., Béthermin, M., et al. 2020, A&A, 633, A90
- Greene et al. (2022) Greene, J., Bezanson, R., Ouchi, M., Silverman, J., & Group, t. P. G. E. W. 2022, arXiv eprints [arXiv:2206.14908]
- Guo et al. (2023a) Guo, Y., Bacon, R., Bouché, N. F., et al. 2023a, Nature, 624, 53
- Guo et al. (2023b) Guo, Y., Bacon, R., Wisotzki, L., et al. 2023b, arXiv eprints [arXiv:2309.06311]
- Guo et al. (2024) Guo, Y., Bacon, R., Wisotzki, L., et al. 2024, Astronomy and Astrophysics, 688, A37
- Harris et al. (2020) Harris, C. R., Millman, K. J., Van Der Walt, S. J., et al. 2020, Nature, 585, 357
- Herenz et al. (2017) Herenz, E. C., Urrutia, T., Wisotzki, L., et al. 2017, A&A, 606, A12
- Hunter (2007) Hunter, J. D. 2007, Computing in Science & Engineering, 9, 90
- Illingworth et al. (2013) Illingworth, G. D., Magee, D., Oesch, P. A., et al. 2013, ApJS, 209, 6
- Inami et al. (2017) Inami, H., Bacon, R., Brinchmann, J., et al. 2017, A&A, 608, A2
- Jaskot & Oey (2014) Jaskot, A. E. & Oey, M. S. 2014, ApJ, 791, L19
- Johnson et al. (2018) Johnson, S. D., Chen, H.-W., Straka, L. A., et al. 2018, ApJ, 869, L1
- Keerthi Vasan G. et al. (2024) Keerthi Vasan G., C., Jones, T., Shajib, A. J., et al. 2024, arXiv eprints [arXiv:2402.00942]
- Kerutt et al. (2022) Kerutt, J., Wisotzki, L., Verhamme, A., et al. 2022, A&A, 659, A183
- Kramarenko et al. (2024) Kramarenko, I. G., Kerutt, J., Verhamme, A., et al. 2024, MNRAS, 527, 9853
- Kusakabe et al. (2022) Kusakabe, H., Verhamme, A., Blaizot, J., et al. 2022, A&A, 660, A44
- Leclercq et al. (2020) Leclercq, F., Bacon, R., Verhamme, A., et al. 2020, A&A, 635, A82
- Leclercq et al. (2017) Leclercq, F., Bacon, R., Wisotzki, L., et al. 2017, A&A, 608, A8
- Leclercq et al. (2022) Leclercq, F., Verhamme, A., Epinat, B., et al. 2022, A&A, 663, A11
- Lee et al. (2014) Lee, K.-G., Hennawi, J. F., Stark, C., et al. 2014, ApJ, 795, L12
- Lehner et al. (2022) Lehner, N., Kopenhafer, C., O’Meara, J. M., et al. 2022, ApJ, 936, 156
- Lehner et al. (2016) Lehner, N., O’Meara, J. M., Howk, J. C., Prochaska, J. X., & Fumagalli, M. 2016, ApJ, 833, 283
- Lofthouse et al. (2020) Lofthouse, E. K., Fumagalli, M., Fossati, M., et al. 2020, MNRAS, 491, 2057
- Lopez et al. (2018) Lopez, S., Tejos, N., Ledoux, C., et al. 2018, Nature, 554, 493
- Luo et al. (2017) Luo, B., Brandt, W. N., Xue, Y. Q., et al. 2017, ApJS, 228, 2
- Martin et al. (2013) Martin, C. L., Shapley, A. E., Coil, A. L., et al. 2013, ApJ, 770, 41
- Maseda et al. (2023) Maseda, M. V., Lewis, Z., Matthee, J., et al. 2023, ApJ, 956, 11
- Matejek & Simcoe (2012) Matejek, M. S. & Simcoe, R. A. 2012, ApJ, 761, 112
- Mauerhofer et al. (2021) Mauerhofer, V., Verhamme, A., Blaizot, J., et al. 2021, A&A, 646, A80
- Meyer et al. (2024) Meyer, R. A., Oesch, P. A., Giovinazzo, E., et al. 2024, arXiv eprints [arXiv:2405.05111]
- Moffat (1969) Moffat, A. F. J. 1969, Astronomy and Astrophysics, 3, 455
- Momose et al. (2014) Momose, R., Ouchi, M., Nakajima, K., et al. 2014, MNRAS, 442, 110
- Morrissey (2018) Morrissey, P. 2018, ApJ, 864, 93
- Méndez-Hernández et al. (2022) Méndez-Hernández, H., Cassata, P., Ibar, E., et al. 2022, A&A, 666, A56
- Oke & Gunn (1983) Oke, J. B. & Gunn, J. E. 1983, ApJ, 266, 713
- Pessa et al. (2024) Pessa, I., Wisotzki, L., Urrutia, T., et al. 2024, arXiv eprints [arXiv:2408.16067]
@@
- Piqueras et al. (2019) Piqueras, L., Conseil, S., Shepherd, M., et al. 2019, Astronomical Data Analysis Software and Systems XXVI ASP Conference Series, 521, 545, conference Name: Astronomical Data Analysis Software and Systems XXVI ADS
- Planck Collaboration et al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., et al. 2020, A&A, 641, A6
- Prochaska et al. (2011a) Prochaska, J. X., Kasen, D., & Rubin, K. 2011a, ApJ, 734, 24
- Prochaska et al. (2011b) Prochaska, J. X., Weiner, B., Chen, H.-W., Mulchaey, J., & Cooksey, K. 2011b, ApJ, 740, 91
- Péroux & Howk (2020) Péroux, C. & Howk, J. C. 2020, ARA&A, 58, 363
- Rafelski et al. (2015) Rafelski, M., Teplitz, H. I., Gardner, J. P., et al. 2015, ApJ, 150, 31
- Rakic et al. (2012) Rakic, O., Schaye, J., Steidel, C. C., & Rudie, G. C. 2012, The Astrophysical Journal
- Richard et al. (2019) Richard, J., Bacon, R., Blaizot, J., et al. 2019, arXiv eprints [arXiv:1906.01657]
- Richard et al. (2021) Richard, J., Claeyssens, A., Lagattuta, D., et al. 2021, A&A, 646, A83
- Rubin et al. (2018) Rubin, K. H. R., O’Meara, J. M., Cooksey, K. L., et al. 2018, ApJ, 859, 146
- Rubin et al. (2011) Rubin, K. H. R., Prochaska, J. X., Ménard, B., et al. 2011, ApJ, 728, 55
- Schroetter et al. (2021) Schroetter, I., Bouché, N. F., Zabl, J., et al. 2021, MNRAS, 506, 1355
- Shaban et al. (2022) Shaban, A., Bordoloi, R., Chisholm, J., et al. 2022, ApJ, 936, 77
- Shapley et al. (2003) Shapley, A. E., Steidel, C. C., & Pettini, M. 2003,ApJ, 588
- Shull & van Steenberg (1982) Shull, J. M. & van Steenberg, M. 1982, ApJS, 48, 95